Pls help me solve this question
With out using table simplify
Log4(2)+log4(32)log4(0.25)

Found 2 solutions by rothauserc, Alan3354:
Answer by rothauserc(4718)

Log4(2) +log4(32)log4(0.25) = log4(2) * (log4(32)log4(0.25))
:
log4(2) = 4^(x) = 2, so log4(2) = 1/2 because x = 1/2
:
log4(32) = 4^x = 32, so log4(32) = 5/2 because x = 5/2
:
log4(0.25) = 4^x = 0.25, so log4(0.25) = -1 because x = -1
:
*************************************************************
Log4(2) +log4(32)log4(0.25) = log4(2) * (log4(32)log4(0.25)) = (1/2) * (5/2) * -1 = -5/4
*************************************************************
:
------------------------
===============================================
"Solve this one" did not come up, only "Edit this solution."
===============================================
So,
Log4(2) +log4(32)log4(0.25) = log4(2) + (log4(32)log4(0.25)) = (1/2) + (5/2) * -1 = 1/2 - 5/2
= -2

Answer by Alan3354(69443)

Pls help me solve this question
---------
Questions are not solved.
Questions are answered.
----
Equations are solved.
Problems are solved.